# The dawn is quiet II: Gaia XP constraints on the Milky Way’s proto-Galaxy from very metal-poor MDF tails

### v1

@@ 5.3 Comparison with Previous Halo MDF Studies
The slopes inferred from Gaia XP metallicity catalogs are broadly consistent with previous constraints on the Milky Way’s metal-poor tail. In the Pristine survey, the inner-halo MDF between $3.4<\mathrm{[Fe/H]}<-2.5$ follows a single exponential with $k\equiv\Delta\log_{10}N/\Delta\mathrm{[Fe/H]}=1.0\pm 0.1$ for main-sequence turnoff stars, establishing a canonical baseline of $k\simeq 1$ (2020MNRAS.492.4986Y). SkyMapper photometric metallicities for $\sim 10^{5}$ giants yield two regimes: $k=1.5$ for $3.0<\mathrm{[Fe/H]}<-2.3$ and $k\simeq 1.0$ for $2.3<\mathrm{[Fe/H]}<-1.5$, with negligible dependence on Galactic height to $|\mathrm{z}|\approx 5$–7 kpc (2019MNRAS.489.5900D; 2021ApJ...911L..23C). These values bracket the range found here from Gaia XP, confirming that the inner-halo MDF tail typically follows an exponential decline with $k\approx 1$.
@@
Other halo MDF studies, with differing spatial coverages and selection functions, have emphasized the metal-poor tail in the outer halo. The Hamburg/ESO Survey revealed a sharp decline near $\mathrm{[Fe/H]}\approx-3.6$, interpreted as evidence for a critical metallicity $Z_{\mathrm{cr}}\approx 10^{-3.4}Z_{\odot}$ required for low-mass star formation (2007MNRAS.381..647S; 2009A&A...507..817S). SDSS Stripe-82 and BOSS spectroscopy described a multi-modal halo MDF with peaks near 1.7-1.7 and 2.3-2.3, corresponding to inner- and outer-halo components with distinct kinematics (2013ApJ...763...65A; 2014A&A...568A...7A). LAMOST giants at $|\mathrm{z}|>5$ kpc show a similar three-component structure-inner and outer halo plus a heated-disk population (2018ApJ...862..163L). The H3 Survey combined with Gaia data indicates that $\gtrsim 95\%$ of halo giants can be attributed to discrete merger remnants: beyond 15 kpc, the halo is dominated by GSE and Sagittarius debris, while the most metal-poor tail arises from lower-mass progenitors (2020ApJ...901...48N; 2021A&A...651A..79B).
@@
5.4 Comparison with Previous GCE Studies
@@
Classical one-zone models show that the low-metallicity tail of the halo MDF should be exponential, with the power set by either lowering the effective yield (mass of freshly produced metals returned to the ISM per unit mass locked in long-lived stars/remnants) by a factor of 13 compared to the disk or adding outflow proportional to star formation (1976ApJ...209..418H). This simple model successfully predicted the number of metal-poor stars down to $\mathrm{[Fe/H]=-3}$ for early observations (1991AJ....101.1865R; 1996AJ....112..668C). However, 2003A&A...404..211P noted that this simple model over-predicts the number of stars below $\mathrm{[Fe/H]<-4}$ if instantaneous recycling approximation (IRA) is relaxed and found that a non-IRA model with a moderate initial inflow and star-driven outflow recovers the observed exponential tail without significant reduction in the effective yield.
In a hierarchical extension, the halo MDF can be written as the sum of subhalo MDFs with mass-dependent yields (2008A&A...489..525P). In this picture, the low-metallicity bins ($\mathrm{[Fe/H]<-2.0}$) are supplied primarily by numerous low-mass stellar subhaloes ($10^{6-7}~\mathrm{M_{\odot}}$). Independently, a merger-tree GCE model finds that the halo MDF requires strong feedback plus pre-enriched infall (2007MNRAS.381..647S). In their fiducial solution, 60% of the metals that pre-enrich the Galactic medium are ejected by halos with a dark matter mass less than $6\times 10^{9}~\mathrm{M_{\odot}}$. Finally, a two-infall GCE model augmented with star-driven outflow reproduces the halo MDF for with a mass-loading factor of 14 and an early infall timescale less than 0.2 Gyr, implying rapid early enrichment and efficient metal loss (2013A&A...554A.135B). These studies depict an open, rapidly evolving proto-Milky Way regulated by gas flows and hierarchical assembly. An early infall episode coupled to outflows enables rapid enrichment while losing a large amount of metals. The lowest-metallicity stars chiefly arise in numerous low-mass progenitors whose ejecta pre-enriched incoming gas.
@@
In our framework, $k$ traces the early gas regulation: shallow tails ($k\simeq 0.6$) arise in models with large initial reservoirs and little subsequent inflow, whereas steeper tails ($k\simeq 1.0$$1.4$) arise for smaller initial reservoirs plus sustained inflow that builds the gas reservoirs while delaying enrichment. The trade-off between SFE and gas cooling also modulate the mapping from the parameter space to MDF unless one parameter takes an extreme value. Fitting the same exponential form to local Auriga star-particle MDFs ($R<10~\mathrm{kpc}$, $|\mathrm{z}|<3~\mathrm{kpc}$) yields a distribution clustered near $k\simeq 1.25$, consistent with a gas-regulated, infall-driven growth scenario under our framework, which we confirm through the global properties of Auriga galaxies. This scenario is also supported by more detailed examination of the gas accretion history of Auriga galaxies by 2022MNRAS.517..832I. This picture is compatible with classical one-zone interpretations in which the tail slope traces the effective yield. A moderate initial gas reservoir of $10^{8-9}\mathrm{M_{\odot}}$ aligns with early-infall models (2003A&A...404..211P; 2013A&A...554A.135B); cooling timescales need only be moderately short-no extreme suppression of the effective yield is required-and low-to-moderate inflow rates are consistent with pre-enrichment by numerous low-mass progenitors (2007MNRAS.381..647S; 2008A&A...489..525P).
5.5 The Case for a Burstless Proto-Galaxy
@@
Cosmological zoom-in simulations show that Milky Way-mass progenitors at $z\gtrsim 2$ undergo bursty star formation. In the cold-mode regime of 2006MNRAS.368....2D, efficient cooling prevents a stable virial shock for halos with $M_{\rm vir}\lesssim 10^{12}\,{\rm M_{\odot}}$ at $z\gtrsim 2$, so dense, cold inflow drives recurrent bursts. In FIRE, these bursts are followed by energetic winds with mass loading $\eta\sim 10$ that expel a substantial fraction of the ISM and transiently depress the SFR (2015MNRAS.454.2691M). In VINTERGATAN, high-$z$ bursts arise from a turbulent, cold ISM reacting to frequent mergers (2022MNRAS.516.2272S). In FIRE-2, the bursty phase ends as a quasi-static, hot inner CGM develops when the halo crosses $M_{\rm vir}\sim 10^{12}\,{\rm M_{\odot}}$ (2021ApJ...911...88S), accompanied by a rapid ($\lesssim 1\,\mathrm{Gyr}$) transition to time-steady, rotation-supported disk star formation (2023MNRAS.519.2598G). Throughout the proto-galaxy phase, the ISM remains turbulent and regulated by coupled inflow-outflow cycles; feedback-driven outflows throttle the net supply to the nascent disk until the accretion mode changes, but do not fully quench star formation (2025ApJ...990....7S).
This early, burst-dominated mode plausibly overlaps the early epoch of disk formation (2021MNRAS.505..889Y). Kinematic diagnostics indicate that dispersion-dominated (low-circularity) orbits are characteristic of stars born during the proto-galaxy and early disk phase (2023MNRAS.523.6220Y; 2024MNRAS.527.6926M), and stars formed in these early bursts can contribute to observed properties at $z\sim 0$ attributed to merger debris such as GSE (2023MNRAS.521..995R). In our GCE framework, a shallow metal-poor tail ($k\approx 0.6$) arises from a large initial star-formation burst followed by ISM heating and weak subsequent inflow, whereas sustained gas accretion and weaker initial bursts generally yields steeper tails ($k\gtrsim 1$). MDFs inferred from Gaia XP catalogs appear smooth-lacking obvious multi-modal structure like the episodic features modeled by 2025OJAp....8E...7T-but measurement uncertainties can wash out burst-scale signatures, especially since the surviving stellar mass in an individual burst is less than $10^{7}\ \mathrm{M_{\odot}}$. 2024arXiv241022250K estimated the stellar mass of the in situ “Aurora” and GSE to be $\sim 3\times 10^{7}\ \mathrm{M_{\odot}}$ between $3<\mathrm{[Fe/H]<-2.0}$ which would prefer a steeper slope ($k\simeq$ 1.0–1.4) based on Fig. 9 c) and  10.
6 Conclusion
@@
    MDFs from all Gaia XP catalogs, after the recommended quality cuts, are well described by single-slope exponentials over $3<\mathrm{[Fe/H]}<-2.0$, with slopes centered near $k\approx 1$ and an interquartile range of 0.5–2.0 (Figs. 5 and 6). The slope remains largely invariant with Galactocentric radius ($R\lesssim 10$ kpc) and apparent magnitude, indicating minimal sensitivity to spatial incompleteness or luminosity-class selection. Modest inter-catalog differences follow known design choices: APOGEE-anchored label-transfer pipelines yield slightly steeper tails ($k\gtrsim 1.2$) owing to limited VMP training coverage, whereas VMP-optimized or CaHK-augmented models produce flatter tails ($k\lesssim 0.8$). Together, these results support a globally smooth, exponential metal-poor tail characteristic of a chemically well-mixed proto-Galactic population.
@@
    Our grid of one-zone GCE models quantifies the direct link between the MDF tail slope and the proto-Galaxy’s gas-regulation parameters. Shallow tails ($k\simeq 0.6$) occur only when the initial cold gas reservoir exceeds $\sim 10^{9}\,\mathrm{M_{\odot}}$, the SFE is below $10^{-10}\,\mathrm{yr^{-1}}$, and inflow contributes less than $20\%$ of the initial mass over the first Gyr. Such models produce an early starburst that rapidly heats and displaces the ISM, leading to a steep decline in the SFR. Conversely, steeper tails ($k\simeq$1.2–1.4) arise for smaller initial reservoirs ($\lesssim 10^{8}\,\mathrm{M_{\odot}}$) with sustained, low-level inflow and SFE of order $10^{-9}\,\mathrm{yr^{-1}}$, indicative of gradual, infall-driven growth (Figs. 8 and 9). The Gaia XP-based slopes cluster near $k\simeq 1.0\pm 0.2$, implying a proto-Galaxy with a moderate initial gas mass of $10^{8-9}\,\mathrm{M_{\odot}}$, moderate inflow, and non-extreme SFE–consistent with a “quiet” early buildup rather than a strongly burst-dominated or inflow-starved phase. Complementary diagnostics such as the cumulative stellar mass-metallicity relation (Fig. 10) will further refine these constraints.
@@
    Among the GCE models that reproduce the observed $\alpha$/Fe]-rise associated with the Galactic spin-up, the modal metal-poor-tail slope is $k\simeq 1.2$. In the Auriga simulation suite, Milky Way-mass analogs exhibit nearly identical slopes ($k\simeq 1.25$) with little galaxy-to-galaxy scatter (Fig. 12). This further demonstrates the constraining power of the observed $\alpha$/Fe]-rise on the proto-Galaxy. These simulated systems assemble their baryons gradually over the first $\sim$1–2 Gyr without a large, isolated initial starburst, closely resembling our $k\simeq 1.0$ GCE models in which gas reservoirs build slowly under low-to-moderate inflow (Fig. 9 and  11). The concordance between the observed, modeled, and simulated slopes indicates a shared early phase in which a turbulent, feedback-regulated ISM throttles net accretion, naturally producing near-exponential metal-poor tails ($k\approx 1-1.2$) prior to the onset of $\alpha$-enhanced refueling at the start of disk formation.
@@
Our results, derived from the chemical imprint of the gas-regulation history, provide a diagnostic of the proto-Galaxy’s initial conditions that is independent of—and more robust than—kinematic signatures. This distinction is crucial because the kinematic record of the earliest phases can be overwritten by later disruptive events. 2025arXiv250909576O demonstrates that massive radial mergers like the GSE can fundamentally scramble orbital information, making the kinematics of the proto-Galaxy stars and stars kicked up from the disc (the Splash) indistinguishable at low metallicities. Their work shows that stellar kinematics at $z=0$ rarely recover the true timing of the disc’s initial spin-up due to this merger-induced scrambling. In contrast, the MDF is a product of the galaxy’s internal gas supply and enrichment history, which is more resilient to such dynamical shuffling. Therefore, our method based on the MDF slope offers a more direct window into the conditions of the pre-merger, proto-Galactic gas reservoir, providing complementary constraints to purely kinematic approaches.
@@
The Gaia XP metallicity catalogs differ in their training sets, flux calibrations, and CMD masks, and residual systematics-particularly near the Galactic plane-may still affect completeness and flatten the counts at $\mathrm{[Fe/H]}\lesssim-2.5$. Our present one-zone framework, while flexible, assumes constant inflow within two one-Gyr phases and does not yet include explicitly time-variable, merger-driven accretion. Future work should incorporate a hierarchical forward model of the MDF that embeds each catalog’s selection function within a multi-zone GCE framework, allowing inflow, star-formation efficiency, and feedback to vary continuously in time while marginalizing over Type Ia delay-time distributions and uncertain yields. The next generation of datasets-Gaia DR4/DR5, WEAVE, and 4MOST-will provide higher-fidelity kinematics and XP-based $\alpha$/M] and [C/Fe] ratios, enabling cleaner separation of disk contaminants and improved recovery of the faintest metal-poor populations. Together, these advances will allow the exponential VMP tail to serve as a precise, physically interpretable observable linking the Milky Way’s earliest gas supply, enrichment, and star-formation history to the emergence of its first long-lived disk.

### v2

@@ V.4 Comparison with Previous Halo MDF Studies
The slopes inferred from Gaia XP metallicity catalogs are broadly consistent with previous constraints on the Milky Way’s metal-poor tail. In the Pristine survey, the inner-halo MDF between $3.4<\mathrm{[Fe/H]}<-2.5$ follows a single exponential with $k\equiv\Delta\log_{10}N/\Delta\mathrm{[Fe/H]}=1.0\pm 0.1$ for main-sequence turnoff stars, establishing a canonical baseline of $k\simeq 1$ (Youakim et al., 2020). SkyMapper photometric metallicities for $\sim 10^{5}$ giants yield two regimes: $k=1.5$ for $3.0<\mathrm{[Fe/H]}<-2.3$ and $k\simeq 1.0$ for $2.3<\mathrm{[Fe/H]}<-1.5$, with negligible dependence on Galactic height to $|\mathrm{z}|\approx 5$–7 kpc (Da Costa et al., 2019; Chiti et al., 2021). These values bracket the range found here from Gaia XP, confirming that the inner-halo MDF tail typically follows an exponential decline with $k\approx 1$.
@@
Other halo MDF studies, with differing spatial coverages and selection functions, have emphasized the metal-poor tail in the outer halo. The Hamburg/ESO Survey revealed a sharp decline near $\mathrm{[Fe/H]}\approx-3.6$, interpreted as evidence for a critical metallicity $Z_{\mathrm{cr}}\approx 10^{-3.4}Z_{\odot}$ required for low-mass star formation (Salvadori et al., 2007; Schörck et al., 2009). SDSS Stripe-82 and BOSS spectroscopy described a multi-modal halo MDF with peaks near 1.7-1.7 and 2.3-2.3, corresponding to inner- and outer-halo components with distinct kinematics (An et al., 2013; Allende Prieto et al., 2014). LAMOST giants at $|\mathrm{z}|>5$ kpc show a similar three-component structure-inner and outer halo plus a heated-disk population (Liu et al., 2018). The H3 Survey combined with Gaia data indicates that $\gtrsim 95\%$ of halo giants can be attributed to discrete merger remnants: beyond 15 kpc, the halo is dominated by GSE and Sagittarius debris, while the most metal-poor tail arises from lower-mass progenitors (Naidu et al., 2020; Bonifacio et al., 2021).
@@
Lowe et al. (2025) provided an independent check on the metal-poor MDF tail using follow-up spectroscopy of Gaia XP–selected candidates. They selected $\sim$6000 Gaia DR3 candidates from the Zhang et al. (2023) catalog, obtained 2dF+AAOmega spectroscopy, and then measured MDF tails for kinematically defined components (halo, prograde/retrograde disk, and GSE). In their Fig. 16, the fitted tail slopes over $3\leq\mathrm{[Fe/H]}\leq-2$ are close to unity (for the halo they report $\Delta\log N/\Delta\mathrm{[Fe/H]}=1.15\pm 0.08$), consistent with the $k\approx 1$ slopes we found from Gaia XP MDF tails over the same metallicity range. In their appendix A, they compared their spectroscopic metallicities to several Gaia XP-based catalogs and found that the level of agreement is strongly pipeline-dependent, with substantial scatter and clear systematics at the metal-poor end. In particular, they show that XP products by Zhang et al. (2023), Andrae et al. (2023a), and Li et al. (2024) tend to assign systematically higher $\mathrm{[Fe/H]}$ to their most metal-poor stars, plausibly driven by reddening-related parameter degeneracies and limited metal-poor coverage in the training sets. This systematic over-estimation of $\mathrm{[Fe/H]}$ at the metal-poor end would shift intrinsically VMP/EMP stars toward higher inferred metallicities, thereby steepening the observed decline. The true metal-poor MDF tail would therefore be flatter (i.e., have a smaller $k$) than that measured from the affected catalogs.
V.5 Comparison with Previous GCE Studies
@@
Classical one-zone models show that the low-metallicity tail of the halo MDF should be exponential, with the power set by either lowering the effective yield (mass of freshly produced metals returned to the ISM per unit mass locked in long-lived stars/remnants) by a factor of 13 compared to the disk or adding outflow proportional to star formation (Hartwick, 1976). This simple model successfully predicted the number of metal-poor stars down to $\mathrm{[Fe/H]=-3}$ for early observations (Ryan & Norris, 1991; Carney et al., 1996). However, Prantzos (2003) noted that this simple model over-predicts the number of stars below $\mathrm{[Fe/H]<-4}$ if instantaneous recycling approximation (IRA) is relaxed and found that a non-IRA model with a moderate initial inflow and star-driven outflow recovers the observed exponential tail without significant reduction in the effective yield.
In a hierarchical extension, the halo MDF can be written as the sum of subhalo MDFs with mass-dependent yields (Prantzos, 2008). In this picture, the low-metallicity bins ($\mathrm{[Fe/H]<-2.0}$) are supplied primarily by numerous low-mass stellar subhaloes ($10^{6-7}~\mathrm{M_{\odot}}$). Independently, a merger-tree GCE model finds that the halo MDF requires strong feedback plus pre-enriched infall (Salvadori et al., 2007). In their fiducial solution, 60% of the metals that pre-enrich the Galactic medium are ejected by halos with a dark matter mass less than $6\times 10^{9}~\mathrm{M_{\odot}}$. Finally, a two-infall GCE model augmented with star-driven outflow reproduces the halo MDF for with a mass-loading factor of 14 and an early infall timescale less than 0.2 Gyr, implying rapid early enrichment and efficient metal loss (Brusadin et al., 2013). These studies depict an open, rapidly evolving proto-Milky Way regulated by gas flows and hierarchical assembly. An early infall episode coupled to outflows enables rapid enrichment while losing a large amount of metals. The lowest-metallicity stars chiefly arise in numerous low-mass progenitors whose ejecta pre-enriched incoming gas.
@@
In our framework, $k$ traces the early gas regulation: shallow tails ($k\simeq 0.6$) arise in models with large initial reservoirs and little subsequent inflow, whereas steeper tails ($k\simeq 1.0$$1.4$) arise for smaller initial reservoirs plus sustained inflow that builds the gas reservoirs while delaying enrichment. The trade-off between SFE and gas cooling also modulate the mapping from the parameter space to MDF unless one parameter takes an extreme value. Fitting the same exponential form to local Auriga star-particle MDFs ($R<10~\mathrm{kpc}$, $|\mathrm{z}|<3~\mathrm{kpc}$) yields a distribution clustered near $k\simeq 1.25$, consistent with a gas-regulated, infall-driven growth scenario under our framework, which we confirm through the global properties of Auriga galaxies. This scenario is also supported by more detailed examination of the gas accretion history of Auriga galaxies by Iza et al. (2022). This picture is compatible with classical one-zone interpretations in which the tail slope traces the effective yield. A moderate initial gas reservoir of $10^{8-9}\mathrm{M_{\odot}}$ aligns with early-infall models (Prantzos, 2003; Brusadin et al., 2013); cooling timescales need only be moderately short-no extreme suppression of the effective yield is required-and low-to-moderate inflow rates are consistent with pre-enrichment by numerous low-mass progenitors (Salvadori et al., 2007; Prantzos, 2008).
V.6 The Case for a Burstless Proto-Galaxy
@@
Cosmological zoom-in simulations show that Milky Way-mass progenitors at $z\gtrsim 2$ undergo bursty star formation. In the cold-mode regime of Dekel & Birnboim (2006), efficient cooling prevents a stable virial shock for halos with $M_{\rm vir}\lesssim 10^{12}\,{\rm M_{\odot}}$ at $z\gtrsim 2$, so dense, cold inflow drives recurrent bursts. In FIRE, these bursts are followed by energetic winds with mass loading $\eta\sim 10$ that expel a substantial fraction of the ISM and transiently depress the SFR (Muratov et al., 2015). In VINTERGATAN, high-$z$ bursts arise from a turbulent, cold ISM reacting to frequent mergers (Segovia Otero et al., 2022). In FIRE-2, the bursty phase ends as a quasi-static, hot inner CGM develops when the halo crosses $M_{\rm vir}\sim 10^{12}\,{\rm M_{\odot}}$ (Stern et al., 2021), accompanied by a rapid ($\lesssim 1\,\mathrm{Gyr}$) transition to time-steady, rotation-supported disk star formation (Gurvich et al., 2023). Throughout the proto-galaxy phase, the ISM remains turbulent and regulated by coupled inflow-outflow cycles; feedback-driven outflows throttle the net supply to the nascent disk until the accretion mode changes, but do not fully quench star formation (Semenov et al., 2025).
This early, burst-dominated mode plausibly overlaps the early epoch of disk formation (Yu et al., 2021). Kinematic diagnostics indicate that dispersion-dominated (low-circularity) orbits are characteristic of stars born during the proto-galaxy and early disk phase (Yu et al., 2023; McCluskey et al., 2024), and stars formed in these early bursts can contribute to observed properties at $z\sim 0$ attributed to merger debris such as GSE (Rey et al., 2023). In our GCE framework, a shallow metal-poor tail ($k\approx 0.6$) arises from a large initial star-formation burst followed by ISM heating and weak subsequent inflow, whereas sustained gas accretion and weaker initial bursts generally yields steeper tails ($k\gtrsim 1$). MDFs inferred from Gaia XP catalogs appear smooth-lacking obvious multi-modal structure like the episodic features modeled by Ting & Ji (2025)-but measurement uncertainties can wash out burst-scale signatures, especially since the surviving stellar mass in an individual burst is less than $10^{7}\ \mathrm{M_{\odot}}$. Kurbatov et al. (2024) estimated the stellar mass of the in situ “Aurora” and GSE to be $\sim 3\times 10^{7}\ \mathrm{M_{\odot}}$ between $3<\mathrm{[Fe/H]<-2.0}$ which would prefer a steeper slope ($k\simeq$ 1.0–1.4) based on Fig. 10 c) and  11.
VI Conclusion
@@
    MDFs from all Gaia XP catalogs, after the recommended quality cuts, are broadly consistent with single-slope exponentials over $3<\mathrm{[Fe/H]}<-2.0$, with slopes centered near $k\approx 1$ and an interquartile range of 0.6–2.0 (Figs. 5, 6, and 7). The slope remains largely invariant with Galactocentric radius ($R\lesssim 10$ kpc) and with parallax-corrected magnitude, indicating limited qualitative sensitivity to spatial incompleteness or luminosity-class selection. Modest inter-catalog differences follow known design choices: APOGEE-anchored label-transfer pipelines yield slightly steeper tails ($k\gtrsim 1.2$) owing to limited VMP training coverage, whereas VMP-optimized or CaHK-augmented models produce flatter tails ($k\lesssim 0.8$). These catalog-dependent effects plausibly explain much of the observed spread in $k$, but they do not alter the main conclusion that the Gaia XP metallicity catalogs consistently recover a broadly smooth, approximately exponential metal-poor tail.
@@
    Our grid of one-zone GCE models quantifies the direct link between the MDF tail slope and the proto-Galaxy’s gas-regulation parameters. Shallow tails ($k\simeq 0.6$) occur only when the initial cold gas reservoir exceeds $\sim 10^{9}\,\mathrm{M_{\odot}}$, the SFE is below $10^{-10}\,\mathrm{yr^{-1}}$, and inflow contributes less than $20\%$ of the initial mass over the first Gyr. Such models produce an early starburst that rapidly heats and displaces the ISM, leading to a steep decline in the SFR. Conversely, steeper tails ($k\simeq$1.2–1.4) arise for smaller initial reservoirs ($\lesssim 10^{8}\,\mathrm{M_{\odot}}$) with sustained, low-level inflow and SFE of order $10^{-9}\,\mathrm{yr^{-1}}$, indicative of gradual, infall-driven growth (Figs. 9 and 10). The Gaia XP-based slopes cluster near $k\simeq 1.0\pm 0.2$, implying a proto-Galaxy with a moderate initial gas mass of $10^{8-9}\,\mathrm{M_{\odot}}$, moderate inflow, and non-extreme SFE–consistent with a “quiet” early buildup rather than a strongly burst-dominated or inflow-starved phase. Complementary diagnostics such as the cumulative stellar mass-metallicity relation (Fig. 11) will further refine these constraints.
@@
    Among the GCE models that reproduce the observed $\alpha$/Fe]-rise associated with the Galactic spin-up, the modal metal-poor-tail slope is $k\simeq 1.2$. In the Auriga simulation suite, Milky Way-mass analogs exhibit nearly identical slopes ($k\simeq 1.25$) with little galaxy-to-galaxy scatter (Fig. 13). This further demonstrates the constraining power of the observed $\alpha$/Fe]-rise on the proto-Galaxy. These simulated systems assemble their baryons gradually over the first $\sim$1–2 Gyr without a large, isolated initial starburst, closely resembling our $k\simeq 1.0$ GCE models in which gas reservoirs build slowly under low-to-moderate inflow (Fig. 10 and  12). The concordance between the observed, modeled, and simulated slopes indicates a shared early phase in which a turbulent, feedback-regulated ISM throttles net accretion, naturally producing near-exponential metal-poor tails ($k\approx 1-1.2$) prior to the onset of $\alpha$-enhanced refueling at the start of disk formation.
@@
Our results, derived from the chemical imprint of the gas-regulation history, provide a diagnostic of the proto-Galaxy’s initial conditions that is independent of—and more robust than—kinematic signatures. This distinction is crucial because the kinematic record of the earliest phases can be overwritten by later disruptive events. Orkney & Laporte (2025) demonstrates that massive radial mergers like the GSE can fundamentally scramble orbital information, making the kinematics of the proto-Galaxy stars and stars kicked up from the disc (the Splash) indistinguishable at low metallicities. Their work shows that stellar kinematics at $z=0$ rarely recover the true timing of the disc’s initial spin-up due to this merger-induced scrambling. In contrast, the MDF is a product of the galaxy’s internal gas supply and enrichment history, which is more resilient to such dynamical shuffling. Therefore, our method based on the MDF slope offers a more direct window into the conditions of the pre-merger, proto-Galactic gas reservoir, providing complementary constraints to purely kinematic approaches.
@@
The Gaia XP metallicity catalogs differ in their training sets, flux calibrations, and CMD masks, and residual systematics-particularly near the Galactic plane-may still affect completeness and flatten the counts at $\mathrm{[Fe/H]}\lesssim-2.5$. Our present one-zone framework, while flexible, assumes constant inflow within two one-Gyr phases and does not yet include explicitly time-variable, merger-driven accretion. Future work should incorporate a hierarchical forward model of the MDF that embeds each catalog’s selection function within a multi-zone GCE framework, allowing inflow, star-formation efficiency, and feedback to vary continuously in time while marginalizing over Type Ia delay-time distributions and uncertain yields. The next generation of datasets-Gaia DR4/DR5, WEAVE (Dalton et al., 2012; Jin et al., 2024), and 4MOST (de Jong et al., 2019)-will provide higher-fidelity kinematics and XP-based $\alpha$/M] and [C/Fe] ratios, enabling cleaner separation of disk contaminants and improved recovery of the faintest metal-poor populations. Together, these advances will allow the exponential VMP tail to serve as a precise, physically interpretable observable linking the Milky Way’s earliest gas supply, enrichment, and star-formation history to the emergence of its first long-lived disk.